# Eye-tracking-Driven Shared Control for Robotic Arms: Wizard of Oz Studies to Assess Design Choices

### v1

@@ 3.5.2. Data analysis
An evaluation of the Areas of Interest (AoI) and the dwell time is necessary to estimate the need for error minimization methods and to determine constraints for the object recognition algorithms. In combination with this technical evaluation of the system, the perception of the non-disabled participants was analyzed. In the Tobii Pro Lab events, Times of Interest (ToI), gaze movements during object selection, and Areas of Interest (AoI) were defined and labeled on the data. Metrics such as dwell times, which describe the total fixation time of an AoI relative to the ToI, and the number of missed hits were calculated by Tobii Pro Lab and exported for further analysis. Based on these parameters, the technical evaluation was performed.
Insights into the factors mentioned in Section 2.3 and Section 2.1 are gained by interpreting gaze data through Times of Interest (ToI), type of gaze movement, and Areas of Interest (AoI) to distinguish dwell times of AoI hits. These parameters can be used to describe the scene based on visual attention. In order to facilitate the understanding of such terms, definitions according to Holmqvist and Nyström (Holmqvist and Andersson, 2017) and Tobii (Pro, 2014) are presented below.
  • Event: Describes in Tobii Pro Lab the time stamp when an event occurs. In this case it marks the start and end time of a task.
  • Time of Interest (ToI): Interval between two related events
  • Area of Interest (AoI): A polygon corresponding to the outline of an object. The resulting area describes a visual stimulus.
  • AoI hit: Describes the entry of the tracked gaze into an AoI.
  • Dwell Time: Describes the time the gaze remains within an AoI after the initial entry. In this study, the gaze was allowed to exit and enter the AoI. It was the sum of time while the gaze stayed within an AoI.
  • Type of eye movement: In this study, fixations and saccades were of most interest. Fixations refer to the time the eyes remain stable and focused on a specific point or area of interest. It was not distinguished from smooth pursuit, in which the head is moved while the gaze rests on a fixed point in the scene. Saccades are rapid changes in gaze direction. They connect fixations. Other eye movements such as vergence movements and microsaccades were not investigated. For further information on gaze movement we refer to Holmqvist and Nyström (Holmqvist and Andersson, 2017).
@@ 4.1. Online survey
20 out of 71 participants completed the survey. Of the 51 participants who filled out the survey, 36 surveys with no responses (N/A) had to be excluded. If participants completed part of the survey and answered more than page two of seven, the surveys were included in the study. Therefore, 34 complete and partially completed surveys were used in the analysis.
Figure 4. Distribution of responses from online survey participants. Q1 to Q3 were answered by participants with disabilities, Q4 to Q6 by family and friends, and Q7 and Q8 from healthcare professionals.
[…]
This bar chart shows the distribution of responses of online survey respondents. Eight questions are presented that highlight the most interesting findings. People with disabilities were asked the following three questions: First question: Did you have assistance in completing the survey? Responses: fully completed surveys 2 yes, 12 no, 0 not answered. Second question: Do you have access to a robotic arm? Responses: fully completed surveys 1 yes, 13 no, 0 not answered. Third question: Could the system presented be used everyday? Responses: fully completed surveys 9 yes, 2 no, 3 not answered. Family members were asked the following three questions: First question: Do you have access to assistive devices? Responses: fully completed surveys 3 yes, 6 no, 0 not answered. Second question: Are you satisfied with the setup of the system? Responses: fully completed surveys 1 yes, 2 no, 6 not answered. Third question: Would the people you care for be able to use the system? Responses: fully completed surveys 2 yes, 1 no, 6 not answered. Caregivers were asked the following two questions: First question: Are you satisfied with the setup of the system? Responses: fully completed surveys 4 yes, 1 no, 6 not answered. Second question: Would the person you care for be able to use the system? Responses: fully completed surveys 6 yes, 0 no, 5 not answered.
Figure 4 summarizes the results of the survey. 14 participants identified themselves in the group of people with disabilities, eight in family members, and 11 in caregivers. Only once did the respondent not select one of these target groups (”other”). The authors decided to include the participant in the family group because of the affiliation description ”Family but without care share”. The results show that most people with disabilities were able to complete the survey on their own (Q1), leading to the conclusion that they have an assistive device that allows them to use a smart device. This conclusion is based on the participants’ statement that most of them cannot move their hands (10 out of 14).
@@
After viewing the video, participants were asked if they or the people they care for would be able to use the system. More than half of the participants confirmed this (17 out of 32; Q3, Q6, and Q8). Only three participants stated the opposite. The caregivers and family members were also asked if they were satisfied with the setup process (Q5 and Q7). Five stated they were satisfied, while three were not. One reason given was the amount of time it took to set up.
@@
Participants were also asked what tasks they would like to accomplish with the system. The group of people with disabilities gave a total of 14 responses: Eating (3 mentions), drinking (2 mentions), reaching for cutlery, pushing buttons (elevator or electric door), applying makeup or brushing their hair, cleaning, scratching, opening a bottle, and lifting objects of various weights from the floor (2 mentions). Family members and caregivers were asked this question twice. First, they were asked to identify situations in which a robotic system could assist with everyday tasks while they were unaware of the system demonstration. Second, they were asked about the presented system. There were 13 responses to the first question. They rated the following tasks as useful: Making the bed, brushing teeth, assisting with eating, thromboses prevention, putting on socks, documentation, going to the bathroom, doing laundry, cleaning (2 mentions), assisting with walking, assisting with doctor’s appointments and correspondence, silent alarm for service calls, and making phone calls. There were 14 responses to the second question: Eating (5 mentions), drinking, turning on lights, operating the aforementioned lift system, brushing teeth, preparing a drink, opening doors, lifting objects (2 mentions), making phone calls and moving objects. In comparison both questions show a high potential for robotic systems and digitization. In the first case, the focus is on assistance with household tasks and communication, while in the second, user-centered tasks are preferred.
@@ 4.2.1. Gaze behavior analysis
The analyses of dwell times in relation to object selection events was used to find technical requirements on the eye-tracking controller. Assumptions such as correlations between object size and dwell time duration were investigated for the dwell time trigger anticipated in the design.
Table 1 shows the distribution of participants’ choice of object selection and information about dwell time. A visual attention map is shown in Figure 5. The chosen interval was 1043 ms before the robot started to move and was based on the time steps made by the evaluation software. Both the data in Table 1 and in Figure 5 used the set intervals. In this time period, events such as searching for the object or deciding which object to select were excluded. It was ensured that the participant only focused on the selected object. The average duration of the test was 2 min 41 sec (SD = 9.8 sec).
The average dwell time was calculated to determine the maximum dwell time limit for the controller, including the time from fixation start until the robot moved. The longest possible dwell time for the participant’s gaze in this setting was found to be 5.612 sec. In this interval, the gaze did not leave the AoI. It indicates that dwell time triggers can be set to high durations, but in the following limitations of the maximum dwell time are presented.
In this study, we assumed that the dwell time is longer the larger the object, due to the larger AoI. Accuracy errors would have less weight than in the small AoIs. The correlation between object sizes and corresponding dwell times were measured and are given in Figure 6. The bottle has a higher standard deviation in dwell time than the cup, implying longer dwell times in some participants’ trials. As seen in Table 1 this is also visible in the dwell times above 500 ms. Reasons why the correlation may not apply are, for example, the fixation of the bottle cap, which is closer to the edge of the AoI than the cup handle. The gaze is more likely to leave the AoI, resulting in lower dwell times and indicating the influence of intention. In comparison, the fork and the block have approximately the same average dwell time and SD, despite having different object sizes. Our first hypothesis was extended that the geometry of the object also influences the dwell time. This influence is indicated by the small width of the fork handle, being less than 1 cm. When accuracy errors occur, it is more likely that gaze is tracked outside the AoI.
@@
The figure shows the experimental setup shown in Figure 1. It is overlaid with a heat map visualizing the attention of all participants during task selection (absolute duration). The heat map is visible on the objects and, in some cases, the area around the objects. Task 1: Three hotspots of low and medium intensity are shown on the light switch. There are few fixations outside the light switch towards the center of the experimental setup. Task 2 - Pick and place: The block shows a high intensity hotspot at the top of the block in the left corner toward the center of the experimental setup. The hotspot extends beyond the boundary of the object. A few fixations to the right of the block are visible in the vicinity of the block. The platform shows an elliptical hotspot perpendicular to the height (y-axis) of the figure with low, medium, and high intensity. Few fixations were found outside the platform above the object. Task 3 - Filling a cup: The cup’s high-intensity hotspot is centered on the top part of the handle and extends slightly beyond the object’s outline. The bottle has a vertically elongated hotspot that starts at a third of the height of the bottle from the bottom and continues to the bottle cap, slightly exceeding the objects outlines. Although only low intensities are shown, the highest intensity is at the bottle cap. Task 4 - Picking a food item: The hotspot on the food item is off-center from the food item itself and above the food item. The hotspot is about the same size as the food item and covers approximately half of the food item itself. A second low-intensity hotspot is shown to the right of the food item, between the food item and the fork. The fork has two hotspots at the base of the fork tines and slightly below in 1/6 of the fork handle. Both hotspots extend across the width of the fork.
@@
The influence of gaze direction was found by examining the low dwell times of the food item placed closest to the user. Looking at the gaze distribution in relation to the scene camera coordinate system, as shown in Figure 7, it is evident that participants tended to move their head to center their gaze when looking at a peripheral object to the right or left. In contrast, participants tended not to move their head when looking down. The detected gaze location was less accurate, due to higher scatter closer to the eye-tracker’s periphery. Physiological impact cannot be excluded such as shadowed pupils by the eyelashes. This resulted in seven missed AoI hits for the food item.
Significant correlation was investigated between dwell time, object size, and gaze direction towards the object. The Shapiro-Wilk test (H = 0.771, df = 97, p = ¡0.001) and the Levene test (H = 2.658, df1 = 6, df2 = 90, p = 0.02) showed no normal distribution or homogeneity. A Kruskal-Wallis test was applied and showed significant differences ($H(6)=17.442$, $p<0.008$). Post-hoc analysis using Dunn’s test with Bonferroni correction revealed significant differences ($p<0.05$) between the following pairs: food item and bottle, food item and platform, food item and light switch, food item and cup, and block and light switch. These results indicate that the measured data for the food item are significantly different from the other objects. Spearman’s $\rho$ was used to analyze the correlation between object size, object position, and dwell time. The frame in Figure 7 was evenly divided into five-by-five sectors in horizontal and vertical directions, so that gaze location could be described by a coordinate map. The analysis showed that dwell time and object size ($\rho=0.373$, $p<0.001$) as well as dwell time and offset in y-direction ($\rho=0.263$, $p=0.009$) were positively correlated. The correlation between the offset values in y-direction and x-direction is weaker and does not correlate with dwell time ($\rho=0.160$, $p=0.116$). This indicates that the assumption that the dwell time is influenced by gaze location towards the object and the object size is correct.
@@ 4.2.2. Perception of hands-on application
In general, satisfaction with the system was expressed in the hands-on study. Participants responded with following comments: 9 positive, 3 negative, 9 times positive and negative, and 3 times with no comments. All statements can be found in Appendix A. Differences were found between the responses of people who work regularly with robotic arms (called prior-experience group below) and people who have no experience with robotic arms (called non-prior-experience group). The non-prior-experience group gave mostly positive comments and focused more on the description of tasks, while the prior-experience group focused on robot kinematics, robot behavior, and user safety, which partially included negative aspects.
While the non-expert group reported being interested, entertained, satisfied, and impressed, they also perceived confusion and uncertainty. The task design itself caused confusion, as participants were not sure how to react to the glass presented by the robot. Considering a real-life application of filling the glass, one of the participants expressed fear of overfilling the glass. Regarding the robotic system, they were satisfied with good reaction time and intuitive control. The velocity was ambivalent, with some participants saying it was just right and others saying it was too slow. Importantly, users reported feeling safer because the robot’s movements were limited to the edge of the table. The fork movement may have had an unsettling effect on the person, as they reported that the fork was not close enough to eat or at the right height for their mouth.
The prior-experience group comments were mostly negative about the robot’s trajectory and the measurement setup. For example, the gripper could crush someone’s hand if the user reached into the setup, due to the lack of pressure sensors. The placement of the emergency stop was a concern for the well-being of future users due to its distance from the participant. The shaky movements of the robot, the non-optimized trajectories, and in one case the robot accidentally touching the bottle while moving to the next object were mentioned. The perceived motor vibration was uncomfortable and the sound of the dropped fork was too loud. Reaction time and object recognition were mentioned as positive. In addition, two participants from both groups stated that they felt uncertain that the robot would stop the task if they looked away or at other objects while the robot was performing the task.

### v2

@@ 3.4.2. Data analysis
Data were recorded using wearable eye-tracking glasses (Tobii, Tobii Pro Glasses 3, accuracy = 0.6). Gaze visualization, data analysis, and post-processing were performed using Tobii Pro Lab (Pro, 2014). In the hands-on study, the computer controlling the robot and visualizing the participant’s gaze was placed 2 m behind the participant so that the user was unaware of the experimenter’s role. Evaluating the areas of interest (AoI) and dwell time is necessary to estimate the need for error minimization methods and determine constraints for object recognition algorithms. Along with this technical evaluation of the system, we analyzed the perception of the non-disabled participants. In the Tobii Pro Lab events, Times of Interest (ToI), gaze movements during object selection, and Areas of Interest (AoI) were defined and labeled on the data. Tobii Pro Lab calculated metrics such as dwell times, which describe the total fixation time of an AoI relative to the ToI and the number of missed hits, and exported them for further analysis. The technical evaluation was performed based on these parameters.
Interpreting gaze data provides insights into the factors mentioned in Section 2.1. ToI, AoI, and dwell time can be used to describe the scene based on visual attention. In order to facilitate the understanding of such terms, definitions according to Holmqvist and Nyström (Holmqvist and Andersson, 2017) and Tobii (Pro, 2014) are presented below.
  • Event: In Tobii Pro Lab, describes the timestamp when an event occurs. In this case, it marks the start and end times of a task.
  • Time of Interest (ToI): The interval between two related events.
  • Area of Interest (AoI): A polygon that corresponds to the outline of an object. The resulting area describes a visual stimulus.
  • AoI hit: The entry of the tracked gaze into an AoI.
  • Dwell Time: The time that the gaze remains within an AoI after the initial entry. In this study, the gaze was allowed to exit and re-enter the AoI. Dwell time is the sum of time while the gaze stays within an AoI.
  • Type of eye movement: In this study, fixations and saccades were of primary interest. Fixations refer to the time that the eyes remain stable and focused on a specific point or area of interest. They were not distinguished from smooth pursuit, which occurs when the head moves while the gaze remains fixed on a point in the scene. Saccades are rapid changes in gaze direction. They connect fixations. Other eye movements, such as vergence movements and microsaccades, were not investigated. Further information on gaze movement can be found in Holmqvist and Nyström (Holmqvist and Andersson, 2017).
@@ 4.1. Online survey
Of the 71 participants, 36 surveys with no responses (N/A) had to be excluded. 19 participants completed the survey, and 15 partially completed surveys were included when they answered at least Q1 to Q12 (referring to the question indicators stated in Appendix A), due to their given insights into daily life. Therefore, the analysis used 34 complete and partially completed surveys. We assume that the large number of empty surveys resulted from people skipping through the survey out of curiosity. Additionally, some institutions reported that they conducted an internal ethical evaluation before distributing the survey to patients and staff. These surveys were automatically saved, even if they were not finished.
The group distribution of the completed surveys was 10 PSMD, 3 family members or friends (fam./fr.), and 6 medical professionals (med. prof.). In comparison, the in the incomplete surveys 4 PSMD, 6 fam./fr., and 5 medical professionals participated. The demographics can be found in Table 1, which summarizes the results of Q1 to Q3. The general and system-related questions varied by group to identify different viewpoints on the system as shown in Figure 3. Table 2 shows the distribution of questions related to the corresponding topic. The results of Q4 to Q15 will be presented in the following topics.
4.1.1. Support situation
For this topic, questions were asked to evaluate the participants’ personal situations and whether there were any obstacles to filling out the survey. Only two of the PSMD group received assistance filling out the survey (Q4(PSMD)). The reason was not stated. Due to limitations in hand and upper body movement, and a lack of access to suitable assistive devices, it can be assumed that the assistance was necessary (Q8, Q9). Relatives and friends only stated once that the person they cared for filled out the survey. Therefore, the risk of bias due to conferring between participants is low.
4.1.2. Level of disability and assistive devices
Q6 revealed that two participants experienced variability in their disability-related limitations from day to day, underscoring the importance of considering individuals with episodic disabilities. These individuals may particularly benefit from multimodal control approaches that can adapt to varying abilities, as discussed by Bhattacharjee et al. and Nanavati et al. (Bhattacharjee et al., 2020; Nanavati et al., 2024b). Regarding Q8, participants differed greatly in their motor abilities and, in most cases, their level of disability remained unchanged (Q7: 12 out of 15). In the other three cases, the level of disability improved, which could indicate stroke rehabilitation. This emphasizes that, for this group, the presented control method may only be applicable for a short time before switching to other control inputs, such as a joystick. This emphasizes the importance of systems that are easy to learn and apply, assisting in short recovery times. Relatives and friends stated that three participants had access to assistive devices, in the form of a patient lift and a repositioning aid.
4.1.3. Robotic arm use
Of all the participants asked, only one stated that she had already used a robotic arm controlled by a chin switch. In the context of Q7(PSMD) and Q8(PSMD), it may be that the level of disability of the other participants does not qualify them to receive a robotic arm from healthcare providers, or they may receive assistance from caregivers or family, which is more common in Germany, Switzerland, and Austria. Approval for a medical device by healthcare providers in Germany can take several months or, in rare cases, years. This indicates that the use of robots is uncommon and can greatly bias the participants’ perception of the system due to its novelty.
4.1.4. WoZ satisfaction
@@
After viewing the video, the participants were asked if they or the person they cared for would be able to use the system. More than half of the participants confirmed this (17 out of 34; Q13 - PSMD, Q14 - Rel./Fr. and Med. Prof.). Only three participants stated the opposite. No reasons were given. Medical professionals and family members were also asked if they were satisfied with the setup process (Q11 - Rel./Fr. and Med. Prof.). Five stated they were satisfied, while three were not. One reason given was the amount of time it took to set up. It might also be influenced by the desire for more information about the system. As Figure 4 shows, the high number of N/A answers in the partially completed surveys occurred after the video. This may be because participants were unable to return to the website after watching the video. Due to missing feedback, this is only a presumption.
Figure 4. Distribution of responses from online survey participants. Q4, Q9 and Q13 were answered by PSMD, Q6, Q11 and Q14 by family and friends and medical professionals.
This bar chart shows the distribution of responses of online survey respondents. Eight questions are presented that highlight the most interesting findings. PSMD were asked the following three questions: First question: Did you have assistance in completing the survey? Responses: fully completed surveys 2 yes, 12 no, 0 not answered. Second question: Do you have access to a robotic arm? Responses: fully completed surveys 1 yes, 13 no, 0 not answered. Third question: Could the system presented be used everyday? Responses: fully completed surveys 9 yes, 2 no, 3 not answered. Family members were asked the following three questions: First question: Do you have access to assistive devices? Responses: fully completed surveys 3 yes, 6 no, 0 not answered. Second question: Are you satisfied with the setup of the system? Responses: fully completed surveys 1 yes, 2 no, 6 not answered. Third question: Would the people you care for be able to use the system? Responses: fully completed surveys 2 yes, 1 no, 6 not answered. Medical professionals were asked the following two questions: First question: Are you satisfied with the setup of the system? Responses: fully completed surveys 4 yes, 1 no, 6 not answered. Second question: Would the person you care for be able to use the system? Responses: fully completed surveys 6 yes, 0 no, 5 not answered.
4.1.5. Task preference
@@
Participants were asked what tasks they would prioritize to accomplish with the system. The group of PSMD gave a total of 14 responses: Eating (3 mentions), drinking (2 mentions), reaching for cutlery, pushing buttons (elevator or electric door), applying makeup or brushing their hair, cleaning, scratching, opening a bottle, and lifting objects of various weights from the floor (2 mentions). Family members and medical professionals were asked this question twice. First, they were asked to identify situations in which a robotic system could assist with daily tasks, while they were unaware of the system demonstration. Second, they were asked about the presented system. Thirteen people responded to the first question. They rated the following tasks as useful: Making the bed, brushing teeth, assisting with eating, thrombosis prevention, putting on socks, documentation, going to the bathroom, doing laundry, cleaning (2 mentions), assisting with walking, assisting with doctor’s appointments and correspondence, silent alarm for service calls, and making phone calls. Fourteen responses were given to the second question. Eating (5 mentions), drinking, turning on lights, operating the aforementioned lift system, brushing teeth, preparing a drink, opening doors, lifting objects (2 mentions), making phone calls, and moving objects. Both questions demonstrate a high potential for robotic systems and digitization. The first focuses on assistance with household tasks and communication, while the second focuses on user-centered tasks. Figure 5 shows the distribution of task preferences mentioned by the participants.
This pie chart shows the classification of all mentioned tasks in relation to the presented system. 28 tasks were grouped into six categories containing the tasks mentioned by the participants and are shown in percentage. The largest category is nutrition with 52%, the second largest category is daily tasks with 19%, the third is mobility with 11%, the fourth is two categories with 7% each described by the terms personal hygiene and well-being. Finally, tasks classified as communication are mentioned with 4%.
@@ 4.2.1. Gaze behavior analysis
Analyses of dwell times in relation to object selection events were used to determine the technical requirements for the eye-tracking controller. Assumptions were investigated, such as the correlation between object size and dwell time duration.
Table 3 shows the distribution of participants’ object selection choices and dwell time information. A visual attention map is shown in Figure 6. The chosen interval was 1,043 ms before the robot started moving, based on the time steps made by the evaluation software. Both data in Table 3 and in Figure 6 represent the intervals. During this time period, events such as searching for the object or deciding which object to select were excluded. Participants were instructed to focus solely on the selected object, to investigate whether natural gaze behavior exhibits gaze cues that hinder implementation or result in insufficient accuracy. The average duration of the test was 2 min, 41 sec (SD = 9.8 sec).
The average dwell time was calculated to determine the maximum limit for the controller’s dwell time, including the time from the start of fixation until the robot moved. In this setting, the longest possible dwell time for the participant’s gaze was found to be 5.612 seconds. During this time, the gaze did not leave the area of interest (AoI). This indicates that dwell time triggers can be set to long durations. However, the following limitations show that such long dwell times are rare.
In this study, we assumed that dwell time would be longer for larger objects due to larger AoIs. Accuracy errors would have less impact on AoI interactions than in the small AoIs. The correlation between object sizes and corresponding dwell times was measured and is given in Figure 7. The dwell time of the bottle has a higher standard deviation than the cup, implying more occurrences of longer dwell times. As seen in Table 3, this is also visible in the dwell times above 500 ms. Reasons why the correlation may not apply are the fixation of the bottle cap, which is closer to the edge of the AoI than the cup handle. The gaze is more likely to leave the AoI, resulting in lower dwell times, which indicates the influence of user intention. In comparison, the fork and the block have similar average dwell times and standard deviations, despite their different sizes. We extended our initial hypothesis to include the influence of geometry of the object on the dwell time. We tested this by evaluating missed hits and gaze behavior with a fork handle that was less than 1 cm wide. When accuracy errors occur, it is more likely that gaze is tracked outside the fork’s AoI.
@@
Examining the low dwell times of the food item placed closest to the user revealed the influence of gaze direction on the dwell time. Looking at the gaze distribution in relation to the scene camera’s coordinate system reveals that participants tended to move their heads to center their gaze when looking at objects to the right or left of them, as shown in Figure 8. In contrast, participants tended not to move their heads when looking down. The detected gaze location was less accurate due to the greater scatter closer to the eye-tracker’s periphery. Physiological factors, such as shadowed pupils by the eyelashes, cannot be excluded to impact dwell time. This resulted in seven missed AoI hits for the food item.
Significant correlations were investigated between dwell time, object size, and gaze direction towards the object. The Shapiro-Wilk test (H = 0.771, df = 97, p = ¡0.001) and the Levene test (H = 2.658, df1 = 6, df2 = 90, p = 0.02) showed no normal distribution or homogeneity. A Kruskal-Wallis test was applied and showed significant differences ($H(6)=17.442$, $p<0.008$). Post-hoc analysis using Dunn’s test with Bonferroni correction revealed significant differences ($p<0.05$) between the following pairs: food item and bottle, food item and platform, food item and light switch, food item and cup, and block and light switch. These results indicate that the measured data for the food item are significantly different from the other objects. Spearman’s $\rho$ was used to analyze the correlation between object size, object position, and dwell time. In Figure 8, gaze location was evaluated by applying a coordinate map visualized by the equally distributed sectors. The analysis showed that dwell time and object size ($\rho=0.373$, $p<0.001$) were positively correlated, as well as dwell time and offset in y-direction ($\rho=0.263$, $p=0.009$). The correlation between the offset values in y-direction and x-direction is weaker and does not correlate with dwell time ($\rho=0.160$, $p=0.116$). This indicates that the assumption that the dwell time is influenced by gaze location towards the object and the object size is correct.
@@ 4.2.2. Perception of hands-on application
In general, the hands-on study revealed satisfaction with the system. In the unstructured interview, the participants’ comments were evaluated based on whether they were positive or negative regarding the robot’s performance. The results were as follows: nine positive comments, three negative comments, nine comments that were both positive and negative, and three comments that were neither positive nor negative. All statements can be found in Appendix B. Differences were found between the responses of people who regularly work with robotic arms (the prior-experience group) and those who have no experience with robotic arms and robots in general (the non-prior-experience group). The non-prior-experience group gave mostly positive comments and focused on describing the tasks, while the prior-experience group focused on robot kinematics, robot behavior, and user safety.
The non-expert group reported feeling interested, entertained, satisfied, and impressed. At the same time, they reported feeling confused and uncertain. The task design itself caused confusion because participants were unsure how to react to the glass presented by the robot. Considering a real-life application of pouring a drink, one participant expressed fear of overfilling the glass. Regarding the robotic system, participants were satisfied with good reaction time and intuitive control. Negative comments concerned the velocity, which created ambivalent reactions. Some participants stated the velocity was just right, and others saying it was too slow. Importantly, users reported feeling safer because the robot’s movements were limited to the edge of the table. The fork movement toward the participant had an unsettling effect, as they reported that the fork was not close enough to eat, or at the right height for their mouth.
Comments from the prior-experience group were mostly negative about the robot’s trajectory and the measurement setup. For instance, the gripper could potentially crush someone’s hand if the user reached into the setup due to the lack of pressure sensors. The placement of the emergency stop was a concern for the well-being of future users, due to its distance from the participant. They also mentioned the robot’s shaky movements, non-optimized trajectories, and one instance where the robot accidentally touched the bottle while moving to the next object. The perceived motor vibration was uncomfortable and the sound of the dropped fork was too loud. Reaction time and object recognition were mentioned as positive, implying that the Wizard of Oz design was convincing.
Additionally, two participants from both groups stated that they were uncertain whether the robot would stop the task if they looked away or at other objects while the robot was performing it, a concern that must be addressed in the control design.